OpenStudy (anonymous):

After the first two term in the sequence below, each odd-number term can be found by dividing the previous term by 2. For example, the 3rd term is equal to 4/2 = 2. Each even-numbered term can be found by adding 8 to the previous term. For example, the 4th term is equal to 2 + 8=10. How many term are there before the first non- integer term? 6,4,... A) 3 B) 4 C) 5 D) 6 E) 7

OpenStudy (anonymous):

E) 7, the sequence continues 6,4,10... then 10/2= 5, 5+8=13, 13/5= 6.5 is not an integer. The 7th term is the first non integer term

OpenStudy (dumbcow):

6,4,2,10 .. 10/2 = 5 5+8 = 13 6,4,2,10,5,13,X 13/2 = 6.5 7th term is a non-integer

OpenStudy (anonymous):

6,4,2,10*